# Joint Learning for Scattered Point Cloud Understanding with Hierarchical Self-Distillation

### v1

@@ IV Experiment
Dataset. Our data is based on three commonly used benchmarks, namely ModelNet40 [41], ScanObjectNN [42], and ShapeNetPart [43]. ModelNet40 is a synthetic mesh dataset consisting of 9,843 training and 2,468 testing samples, categorized into 40 entities. On the other hand, ScanObjectNN is a challenging real-world dataset with 15,000 objects classified into 15 categories, comprising 2,902 unique object instances. To generate scattered point clouds, instead of randomly masking out points, we sparsely sample a small number of points to be patch centroids using the farthest point sampling (FPS) algorithm, and subsequently attach the surrounding points using the kNN algorithm. Especially, we have three data types as input for ModelNet40: with the size of , , and (the 8 neighbors are sampled based on a dense resolution of the ground truth (GT), making it much sparser than the previous two types), respectively. For ScanObjectNN and ShapeNetPart, we randomly sample points as input.
@@
Data Preparation. We sample the scattered point clouds based on specific number of seeds and neighbors based upon FPS and kNN. As such, we list the sparsity types for all datasets we have used in Table I. Moreover, in Fig 4, we visualize the Sparsity for different types of inputs for experiments on ModelNet40. Specifically, the input with points has the most scattered distribution as the neighbors are sampled from a dense version, i.e., 10,000 points. The data pre-processing code for all datasets is included in the source code.
Implementation Detail. Our networks are trained on a server deployed with 10 Nvidia RTX 2080Ti GPUs, with a batch size of 260 for classification and 160 for segmentation during training and testing for all models. We do not apply the voting strategy throughout experiments. Except otherwise noted, we omit the upstream name (i.e., SnowflakeNet) and directly use the downstream model names to denote two-stage methods, and use for our joint learning framework.
Evaluation Metrics. Following the standard in the literature, we report the overall accuracy (OA), mean-class accuracy (mAcc), and Chamfer distance (CD) () for classification, and mean-instance and mean-class intersection over union (mIOU/cIOU) for segmentation.
IV-A Classification with Self-Distillation
@@
In this study, we investigate the performance of PointNet++, CurveNet, and PointMLP as classification-only models. To demonstrate the difference between these methods and our proposed HSD framework, we evaluate their accuracy on all downstream networks with full input, as presented in Table II, indicating that HSD is capable of refining the classification results. Since the input point cloud is complete, we only take one view of it for this task. To ensure a fair comparison, we re-train all baselines on the same machine as the proposed models. The data is exactly same as the one used in other canonical models: we take 1,024 points as input and the training/testing splitting follows the standard criterion in the community. The best testing results of OA and mAcc are reported. In this task, we utilize the cosine annealing scheduler to establish the learning rate, as same as PointMLP. The minimum learning rate is 0.005.
IV-B Qualitative Visualization
Figs. 5, 6, and 7 visually demonstrate the improvements in the quality of the reconstructed point clouds for ModelNet40, ScanObjectNN, and ShapeNetPart, respectively. To enhance visualization, we employ the PB_T50_RS without background for ScanObjectNN in Fig. 6. Notably, some input objects are unrecognizable even for human, highlighting the necessity of the upstream stage for better understanding of 3D objects. Particularly, in Fig. 5, the chair’s leg can be essential to the final classification, and our HSD methods successfully reconstruct this part. In Fig 7, the quality of reconstructed point clouds is even better than the insufficiently sampled GT. Since the final objective is object recognition rather than completion, the reconstructed point clouds are restricted to have 1,024 points to ensure downstream efficiency, which generally recovers only the overall shape of an incomplete point cloud.
@@ IV-B Qualitative Visualization
IV-C Incomplete Point Cloud Classification
The point sets collected by real sensors (e.g., Lidar) are only distributed on the object’s surface and only near the side of Lidar, resulting in misalignment of classification and uncertainty of boundary detection. The upstream network tackles this challenge by reconstructing the incomplete point cloud into a complete shape. In this task, all of our models utilize incomplete point clouds as input. All two-stage models are built with full architectures, while our downstreams are elite versions. The size of the target point cloud is 1,024 for fair comparisons. During training, there are 8 views for each incomplete input. In this task, we gradually decay the learning rate for every 50 epochs with a factor of 0.8. The initial learning rate is 0.001. In following experiments, we denote as our joint learning methods.
Results on ModelNet40. The results in Table III reflect that the proposed HSD models outperform the baselines in terms of OA for incomplete object classification. We also conduct completion and classification tasks separately, rather than in an end-to-end scheme. The results generated by SnowflakeNet are directly fed into the fully trained downstream classification networks. For PointNet++, due to unstable predicted results, we run the evaluation for 4 times and report the average for both OA and mAcc. Our HSD models achieve the best OA with competitive CD compared to the two-stage methods. It can be also observed that our proposed methods significantly outperform the two-stage methods on the sparsely masked data. We infer that our joint learning models enhance the information gain especially on sparsely distributed data, mimicking the inference procedure in the real scenario. It can be observed that HSD works less effective on CurveNet. Empirically, despite CurveNet has a hierarchical nature, it walks among the points, forming the curves to represent local features. Unlike query ball and kNN in conventional hierarchy-based methods, curves share more discriminative features in the same region, making different heads share less mutual information.
Results on ScanObjectNN. In this task, we follow the standard in the community and use the hardest variant of ScanObjectNN: PB_T50_RS with background. In Table IV, although our baseline models (i.e., without HSD) have already achieved competitive performances, our HSD models achieve superior performances in terms of OA, mAcc, and CD, comparing to SOTA methods that are specifically trained on ScanObjectNN with complete shapes. Notably, all of our jointly learned HSD models already outperform the most-recent hierarchy-based methods Point-PN and PointConT. This highlights the effectiveness and importance of the proposed method on the real-world setup. More results on ScanObjectNN without background are given in Table. V. Here, we primarily compare our methods to our baselines, except for the upstream’s reconstruction error without joint learning. It can be again observed that our proposed methods achieve lower CD than merely using a completion network SnowflakeNet.
@@ IV-C Incomplete Point Cloud Classification
IV-D Quantitative Analysis
@@
Unlike images, one can only define probability distribution for continuously distributed variables, such as point clouds. In this section, we only measure information by discretization and relative information content in terms of KL divergence and cross entropy. See the supplemental for details of discretization. First, we visualize the information plane in Fig. 8(a), where both and gradually increase and eventually hit the maximum. This makes sense, since our MAE’s objective is to fully reconstruct from latent representations . Similar phenomenon also occurs in an autoencoder-based classifier [45]. This is also in line with other findings in [46, 47], where compression is not always necessary for good generalization performance. Second, since Fig. 8(a) merely differentiate the information of 3 levels separately, we quantify in terms of KL loss in Fig. 8(b). The differences ( and ) between companions and the teacher are decreasing, indicating that the information from the deepest level is flowing to previous levels, providing a stronger regularization. Finally, we utilize cross entropy to approximate in Fig. 8(c) , which suggests that the teacher shares most mutual information to real labels.
Specifically, we use 6 bins to digitize the information (i.e., network weights) from the 3 levels for the initial 100 epochs. Thereafter, with the probabilities of the given data and labels , mutual information and is therefore estimated via:
IV-E Ablation Study
@@
Sections IV-A and IV-C present the comparisons between our baselines and the models based on HSD, indicating the superior performance of the proposed HSD mechanism. The results also showcase the extendibility of the HSD mechanism in scenarios involving flawed or rapidly scanned point clouds. To validate the effectiveness of HSD, we conduct ablation studies with or without the self-distillation loss, on the PB_T50_RS variant (without background) of ScanObjectNN. Here, PointNet++ serves as the backbone for the downstream model, and the results are presented in Table VI. First, the performances of baseline and DSN (i.e., w/o. HSD as discussed in section III-D) are reported. In addition, the significance of the teacher’s code length is demonstrated by employing identical dimensions in the hyperspaces for the last two classification heads, as indicated by the same code length (SCL). Lastly, the full setting with HSD further regularizes for the optimization. These two observations suggest that a longer code length provides more information (i.e., ), and maximizing the mutual information provides stronger regularization.
@@
Worth mentioning, the selection of MAE in our method is also plug-and-play, as shown in Table VII, where we use PMP-Net++ [22] as the upstream. Note that PMP-Net++ is simplified analogue to SnowflakeNet. With the same downstream network, i.e., PointNet++-HSD, Our model with a more powerful upstream again improves all three metrics. This verifies again HSD helps MAE to rearrange high-level details that cannot be reconstructed by MAE alone.
IV-F Part Segmentation
@@
To demonstrate the generality of our framework, we compare our methods to the SOTAs of part segmentation in Table VIII. Since PointMLP and PointNeXt use both coordinate and normal information for training, we retrain them with only the coordinates (2,048 points) on the training set of ShapeNetPart, and test the result generated from the upstream. Both training and testing sets are normalized for consistency. Under an unfair condition, i.e., with merely one half of input (1,024 points) for training, our simplest model (PointNet++-HSD) already outperforms PointMLP. This suggests that our framework possesses the potential to elevate the performance of a model to that of a subsequent generation, and reaffirms the generalization ability of our framework for scattered point clouds on the part segmentation task. We elaborate the results of per-class segmentation IOU in Table IX Specifically, the per-class IOUs are obtained from the checkpoints of the best cIOU for fair comparisons.
Implementation of Part Segmentation. The loss computation and evaluation of incomplete point cloud segmentation are non-trivial. The per-point accuracy and mean IOU’s computation necessitate precise point order as in the ground truth data. Given that the generated point cloud lacks this precise order, direct computation of these metrics becomes unfeasible. However, the reconstruction loss presents an opportunity to roughly compute them, as it formulates the bidirectional mapping, albeit not on a per-point basis, between the predicted and ground truth point clouds. Based upon this mapping, the point index orders are aligned between the two point sets.
IV-G Failure Case
ModelNet40 encompasses a substantial number of classes that exhibit similarities in shape, e.g., table and bench, known as similar-class issue. We consider the similar-class issue as a partial failure case, especially when the sparsity level is low or medium. Another potential failure case is for extremely complex shapes, which may not be appropriately recovered.
V Conclusion
@@
Recent state-of-the-art models have shown less effectiveness on incomplete point clouds. In this paper, we have proposed a cascaded network that tailors the upstream and downstream networks for analyses on scattered point clouds. In our method, MAE serves as a rudimentary form of object completion focusing on low-level features, while HSD further refines the understanding of this preliminary reconstruction focusing on high-level features. Our experiments demonstrate that our joint learning design performs good generalization on scattered point clouds, indicating that the reconstructed objects are beneficial for better environmental understanding, particularly in scenarios where perfect scanning is impractical during data collection. Moreover, our proposed hierarchical self-distillation improves the capacity of hierarchy-based classification models by maximizing the mutual information among layers, indicating that the optimization process is further regularized by the information flow.

### v3

@@ 4 Experiment
Dataset. Our data are sourced from three widely recognized benchmarks: ModelNet40 [49], ScanObjectNN [50], and ShapeNetPart [51]. ModelNet40, a synthetic mesh dataset, includes 9,843 training and 2,468 testing samples, classified into 40 categories. In contrast, ScanObjectNN presents a challenging real-world dataset comprising 15,000 objects across 15 categories, including 2,902 unique instances. To generate scattered point clouds, we employ a sampling strategy that captures representative patches: instead of randomly masking out points, we use farthest point sampling (FPS) to select patch centroids, then incorporate surrounding points via the kNN algorithm. For ModelNet40, we prepare three input types with the following sizes: , , and , where the latter configuration uses eight neighbors sampled at a denser resolution from the ground truth (GT), making it much sparser than the other configurations. For ScanObjectNN and ShapeNetPart, we randomly sample points as input.
@@
Data Preparation. We sample scattered point clouds based on a specific number of seeds and neighbors, leveraging FPS and kNN to generate varied sparsity levels. Table 1 details the sparsity configurations used for each dataset, and Fig. 3 illustrates the sparsity variations in inputs for ModelNet40. Specifically, the configuration with points has the most scattered distribution, as neighbors are selected from a dense point set of 10,000 points. This creates a higher sparsity type, where each cluster has limited range coverage yet high local density. Our sparse sampling methodology inherently mirrors random masking but operates additively rather than by reduction. This approach emulates a continuous scanning scenario, simulating an object being progressively scanned from various perspectives with potential occlusions. Unlike random masking, which simply omits regions, this method emphasizes a more focused understanding of single objects, making it well-suited for understanding tasks. The data pre-processing code for all datasets are provided within our source code.
Implementation Detail. Our networks are trained on a server deployed with 10 Nvidia RTX 2080Ti GPUs, using a batch size of 260 for classification tasks and 160 for segmentation tasks across all models, both during training and testing phases. In the downstream, an upsampling ratio of (1, 1, 2) is applied by default to point-splitting modules. For consistency, we omit the upstream network name (i.e., SnowflakeNet) and directly use the downstream model names to denote two-stage methods; we refer to our joint learning framework as . Take our PointNet++-HSD for instance, we configure to gradually increase coverage ranges. To maintain experimental simplicity, we do not apply a voting strategy in any experiments. Our source code is available at: https://github.com/ky-zhou/PointHSD.
Evaluation Metrics. In line with standard practices in the literature, we report the following metrics: Overall accuracy (OA), mean-class accuracy (mAcc), and Chamfer distance (CD) () for classification, and mean-instance and mean-class intersection over union (mIOU/cIOU) for segmentation.
4.1 Classification with Self-Distillation
In this study, we evaluate the classification performance of PointNet++, CurveNet, and PointMLP as baseline models, comparing their accuracy against our proposed Hierarchical Self-Distillation (HSD) framework. To underscore the impact of HSD on classification, we test each downstream network with full point set inputs, as shown in Table 2. The results indicate that HSD enhances classification outcomes by refining feature quality across layers. Since the input point cloud is complete, only a single viewpoint is used for this task. For fair comparison, we re-train all baseline models on the same machine and with identical data splits as our HSD-based models. Following standard community practices, we use 1,024 input points. The highest testing results for OA and mAcc are reported. To regulate the learning rate, we employ a cosine annealing scheduler with a minimum rate of 0.005 as same as PointMLP. It is noteworthy that HSD for complete shapes operates similarly to DSN as the feature quality is typically sufficient after the initial extraction layer. Optimal performance in these cases may arise from intermediate layers, which aligns with the understanding that, for complete objects, accuracy hinges on the selection of local kernel coverage at each layer. This suggests that HSD may be particularly adept at handling scattered point clouds. In the following section, we present quantitative results that demonstrate the effectiveness of HSD applied to sparse data.
4.2 Incomplete Point Cloud Classification
@@
Point sets captured by real sensors, such as LiDAR, are typically distributed only on an object’s surface, predominantly on the side facing the sensor, which often leads to classification misalignment and challenges in boundary detection. The upstream network tackles this challenge by reconstructing the incomplete point cloud into a complete shape. In this task, all of our models process incomplete point clouds as input. All two-stage models are built with full architectures, while our downstreams are elite versions. The target point cloud size is set to 1,024 for fair comparisons. During training, each incomplete input is augmented with 8 different views. Additionally, we apply a gradual learning rate decay by a factor of 0.8 every 50 epochs, starting with an initial learning rate of 0.001. In following experiments, we denote as our joint learning methods.
Results on ModelNet40. The results in Table 3 reflect that the proposed HSD models outperform the baselines in terms of OA for incomplete object classification. We also conduct completion and classification tasks separately rather than in an end-to-end scheme. In particular, the results generated by SnowflakeNet are directly utilized as inputs to fully trained downstream classification networks. For PointNet++, given the variability in predictions, we perform the evaluation four times and report the average results for both OA and mAcc. Our HSD models achieve the highest OA with competitive CD compared to two-stage methods. Notably, our proposed methods significantly outperform two-stage approaches on sparsely masked data. We infer that our joint learning models enhance the information gain especially on sparsely distributed data, mimicking the inference procedure in the real-world scenarios. It is also observed that HSD is less effective when applied with CurveNet.
Results on ScanObjectNN. In this task, we follow the standard in the community and use the hardest variant of ScanObjectNN: PB_T50_RS with background. In Table 4, although our baseline models (i.e., without HSD) have already achieved competitive performances, our HSD models demonstrate superior results in terms of OA, mAcc, and CD, outperforming SOTA methods that are specifically trained on complete shapes from ScanObjectNN. Notably, all of our jointly learned HSD models surpass the most-recent hierarchy-based methods Point-PN [38] and PointConT [41], underscoring the effectiveness and relevance of our approach in real-world scenarios. More results on ScanObjectNN without background are given in Table. 5. Here, we primarily compare our methods to our baselines, except for the upstream’s reconstruction error without joint learning. Once again, our proposed methods achieve lower CD compared to using a standalone completion network like SnowflakeNet.
Why HSD works less effective on CurveNet? Empirically, despite CurveNet has a hierarchical nature, it walks among the points, forming the curves to represent local features. Unlike query ball and kNN commonly used in traditional hierarchical methods, curves share more discriminative features in the same region, resulting in less mutual information shared among different levels.
4.3 Quantitative Analysis
Unlike images, probability distributions can only be defined for continuously distributed variables. For point clouds, we assess information by discretizing and evaluating relative information content using KL divergence and cross-entropy. First, we visualize the information plane in Fig. 4(a), where both and steadily increase, eventually reaching a maximum. This is reasonable, since the objective of our MAE is to fully reconstruct from latent representations . A similar phenomenon is observed in an autoencoder-based classifier [53], and aligns with other studies, such as [54, 55], which indicate that compression is not always required for strong generalization performance. Second, since Fig. 4(a) merely differentiate the information of 3 levels separately, we quantify in terms of KL loss in Fig. 4(b). The differences ( and ) between companion levels and the teacher measured by KL divergence () decrease progressively. This indicates that information from the deepest level flows to preceding levels, providing them with stronger regularization. The observation that further supports the assumption in Sec. 3.4 that for a sparse input. Finally, we utilize cross-entropy to approximate in Fig. 4(c), which suggests that the teacher shares the most mutual information to true labels and that information becomes richer as increases.
Specifically, we use 6 bins to discretize the information (i.e., network weights) across the three levels during the initial 100 epochs. With the probabilities of the given data and labels , mutual information and are estimated as follows:
4.4 Qualitative Visualization
Figs. 5, 6, and 7 visually demonstrate the improvements in reconstructed point cloud quality for ModelNet40, ScanObjectNN, and ShapeNetPart, respectively. For enhanced visualization, Fig. 6 employs the PB_T50_RS without background for ScanObjectNN. Notably, some input objects are unrecognizable even for human, highlighting the necessity of the upstream stage for better understanding of 3D objects. Particularly, in Fig. 5, the reconstruction of the chair’s leg—a crucial feature for accurate classification—is successfully achieved with our HSD methods. In Fig. 7, the reconstructed point clouds exhibit higher quality than the sparsely sampled ground truth. Since the final objective is object recognition rather than precise completion, the reconstructed point clouds are restricted to 1,024 points to maintain efficiency in downstream tasks, capturing only the overall shape of an incomplete point cloud.
4.5 Ablation Study
@@
Sections 4.1 and 4.2 present the comparisons between our baseline models and those based on the HSD mechanism, indicating the superior performance of the proposed approach. The results further demonstrate the scalability of the HSD mechanism in scenarios involving flawed or rapidly scanned point clouds. To validate the effectiveness of HSD, we conduct ablation studies with and without the self-distillation loss, on the PB_T50_RS variant (without background) of ScanObjectNN. Here, PointNet++ serves as the backbone for the downstream model, with results presented in Table 6. First, the performances of baseline and DSN (i.e., w/o. HSD as discussed in section 3.4) are reported. In addition, we demonstrate the importance of a suitably chosen teacher’s code length by employing identical dimensions in the hyperspaces for the last two classification heads, as indicated by the same code length (SCL). Lastly, the full HSD setting provides stronger regularization for the optimization. These two observations suggest that a reasonably longer code length to hold the feature extracted by a larger perceptual field provides more information (i.e., ), and maximizing the mutual information leads to stronger regularization. Specifically, compared to DSN, i.e., strong regularization, our HSD reduces reconstruction error by 3% and increases overall accuracy by 0.4%, confirming the effectiveness of the proposed regularization term.
@@
Worth mentioning, the selection of MAE in our method is also plug-and-play, as shown in Table 7, where we use PMP-Net++ [22] as the upstream. Note that PMP-Net++ is simplified similarly as SnowflakeNet. With the same downstream network, i.e., PointNet++-HSD, Our model with a more powerful upstream consistently improves all three metrics. This further verifies that HSD enables MAE to rearrange high-level details that cannot be reconstructed by the MAE alone.
4.6 Part Segmentation
@@
To demonstrate the generality of our framework, we compare our methods with the SOTAs for part segmentation in Table 8. Since PointMLP [13] and PointNeXt [56] use both coordinate and normal information for training, we retrain them with only the coordinates (2,048 points) on the ShapeNetPart training set and test the results generated from the upstream. Both the training and testing sets are normalized for consistency. Under an unfair condition, i.e., with merely half of the input (1,024 points) for training, our simplest model (PointNet++-HSD) already outperforms PointMLP. This suggests that our framework possesses the potential to elevate the performance of a model to that of a subsequent generation, and reaffirms the generalization ability of our framework for scattered point clouds on the part segmentation task. We present the per-class segmentation IOUs in Table 9. Specifically, the per-class IOUs are obtained from the checkpoints of the best cIOU for fair comparisons.
Implementation of Part Segmentation. The network structure for part segmentation closely resembles that of classification, with the key distinction that the multi-heads leverage a class-agnostic shape code for self-distillation. However, Calculating loss and evaluating segmentation on incomplete point clouds are non-trivial. Metrics such as per-point accuracy and mean IoU require precise point ordering, as in the ground truth data. Since the generated point cloud lacks this exact order, directly computing these metrics becomes infeasible. The reconstruction loss offers a workaround by defining a bidirectional mapping—though not on a per-point basis—between the predicted and ground truth point clouds. This mapping enables point index alignment between the two sets.
4.7 Failure Case and Future Work
ModelNet40 contains a diverse set of classes, some of which have similar shapes (e.g., table and bench), which we refer to as the similar-class issue. We consider this a partial failure case, particularly under conditions of low to medium sparsity. Additionally, extremely complex shapes may pose challenges for accurate recovery.
5 Conclusion
@@
Recent state-of-the-art models have shown limited effectiveness on incomplete point clouds. In this paper, we have proposed a cascaded network that tailors the upstream and downstream components for analyzing scattered point clouds. In our method, MAE serves as a rudimentary form of object completion, focusing on low-level features, while HSD further refines the understanding of this preliminary reconstruction by emphasizing high-level features, making it particularly well-suited for scattered point cloud analysis. Our experiments demonstrate that our joint learning design generalizes well to scattered point clouds, indicating that the reconstructed objects enhance environmental understanding, particularly in scenarios where perfect scanning is impractical during data collection. Moreover, our proposed hierarchical self-distillation improves the capacity of hierarchy-based classification models by maximizing the mutual information between layers, thereby further regularizing the optimization process through information flow.